OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (2, − 2) and a directrix of y = 0?

OpenStudy (anonymous):

f(x) = −one fourth (x − 2)^2 − 1 f(x) = −one fourth (x − 2)^2 + 1 f(x) = −one fourth (x − 2)^2 f(x) = one fourth (x − 2)^2 + 1

OpenStudy (anonymous):

@myininaya @amistre64

OpenStudy (anonymous):

@austinL

OpenStudy (anonymous):

@surjithayer

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

vertex will be half way between Focus and directrix Vertex = ?

ganeshie8 (ganeshie8):